OpenStudy (anonymous):

solve the equation 4x+7y=28 for y in terms of x

OpenStudy (anonymous):

4x + 7y = 28 transposing 7y to RHS (right hand side) 4x = 28 - 7y 4x + 28 = -7y 4(x + 7) = -7y 4(x + 7)/-7 = y -4(x+7)/7 = y (-4x + 28)/7 = y (28-4x)/7

OpenStudy (anonymous):

did u understand tell me if yes than give me a medal

OpenStudy (radar):

I didn't understand the transition: 4x=28-7y to 4x+28=-7y shouldn't it of been 4x-28=-7y ????

OpenStudy (anonymous):

radar, you're right

OpenStudy (radar):

I think so. 4(x-7)=-7y 4(x-7)/-7=y (4x-28)/-7=-(4x/7)+4 y=4-(4x/7)